OpenStudy (anonymous):

When we finally figure out how to travel faster than the speed of light and can make a trip to a neighboring star system 4 light years away in six days, what is our average velocity in miles per hour ? 58,697,136,000,000 miles = 10 light years

OpenStudy (anonymous):

\[\frac{58697136000000}{((10*365)+2) 24}=6.69692*10^8 \text{mph} \]

OpenStudy (anonymous):

Sorry. Did not answer the question as presented. The distance to the object star system is (0.4)*(58,697,136,000,000) miles. Six days is 6*24 hours. The answer in miles/hour is:\[\frac{(4/10)*(58697136000000)}{(6)* 24}=163,047,600,000 \text{ mph} \]That's One Hundred Sixty Three billion, 47 Million, Six Hundred Thousand miles per hour.